OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16

OpenStudy (anonymous):

x = -9 x = -1

OpenStudy (texaschic101):

-4 |x + 5 | = -16 -- divide by -4 | x + 5 | = 4 x + 5 = 4 -(x + 5) = 4 x = 4 - 5 -x - 5 = 4 x = -1 -x = 4 + 5 -x = 9 x = -9 I agree with musa :)